?. The following data are oxygen uptake rate of bakers yeast for enzymatic a particular medium at 25? and various oxygen partial pressure were reaction in obtained. Determine kinetic parameters (Kn and V) Po2 (mmHg) 0 0.5 1.0 1.5 2.5 3.5 5.0 Qo2 (M/h) 23.5 33.0 37.5 42.0 43.0 43.0
Answer #1

The initial rate (in the Question denoted as QO2) of an enzyme-catalyzed reaction can be described by Michaelis–Menten equation which is given by:

v_{0}=\frac{v_{max}\,p_{O_2}}{K_m+p_{O_2}}

Using the first three data one can write the initial rates as:

23.5=\frac{v_{max}\times 0.5}{K_m+0.5}                --------------------(1)

33=\frac{v_{max}\times 1}{K_m+1}                      --------------------(2)

Therefore dividing (1) by (2) one have:

\frac{k_m+0.5}{k_m+1}\times \frac{23.5}{33}=0.5

So, km=0.678 mmHg.

Using the value of km in (1) one get:

     v=55.374 M/h